A gas has a volume of 120m³ at stp. if the gas is compressed adiabatically to a volume of 25m³. calculate the new temperature, pressure and work done on the gas. [Cp= 1.68kj/kg-¹ k-¹, Cv = 1.20 kj kg-¹ k-¹]


1
Expert's answer
2021-11-29T19:17:13-0500

For adiabatic compression

"P_{2}=P_{1}(\frac{V_{1}}{V_{2}} )^r"

Where:

"P_{1}=" initial pressure= atmospheric pressure"=1.00\u00d710^5N\/M^{2}"

"P_{2}=" final pressure

"V_{1}=" Initial volume="120M^{3}"

"V_{2}=" final volume="25M^3"

"r=\frac{Cv}{Cp}"


"r=" "\frac{1.68kj/kg^{-1}K^{-1}}{1.2kj/kg^{-1}K^{-1}} =1.4"

Therefore:

"P_{2}=(1×10^5N/M^2)×(\frac{120M^3}{25M^3})^{1.4}=8.99×10^5N/M^2"


From Ideal gas laws Final Temperature:

"T_{2}=(\frac{P_{2}V_{2}}{P_{1}V_{1}})T_{1}"

Where:

"T_{1}= Room temperature=293K"

"T_{2}= (\frac{8.99×10^5N/M^2×25M^3}{1×10^5N/M^2×120M^3})×293K=548.76K"


Work done in Adiabatic process W:

"W=\frac{1}{1-r}(P_{2} V_{2}-P_{1} V_{1})"


"W=\frac{1}{1-1.4}(8.99×10^5N/M^2×25M^3-1×10^5N/M^2×120M^3)= -261.875j"